# Toward Safe and Human-Aligned Game Conversational Recommendation via Multi-Agent Decomposition

### v2

@@ Abstract
Conversational recommender systems (CRS) have advanced with large language models, showing strong results in domains like movies. These domains typically involve fixed content and passive consumption, where user preferences can be matched by genre or theme. In contrast, games present distinct challenges: fast-evolving catalogs, interaction-driven preferences (e.g., skill level, mechanics, hardware), and increased risk of unsafe responses in open-ended conversation. We propose MATCHA, a multi-agent framework for CRS that assigns specialized agents for intent parsing, tool-augmented retrieval, multi-LLM ranking with reflection, explanation, and risk control which enabling finer personalization, long-tail coverage, and stronger safety. Evaluated on real user request dataset, MATCHA outperforms six baselines across eight metrics, improving Hit@5 by 20%, reducing popularity bias by 24%, and achieving 97.9% adversarial defense. Human and virtual-judge evaluations confirm improved explanation quality and user alignment. Code will be released upon acceptance.
@@
Zheng Hui[Uncaptioned image] [Uncaptioned image], Xiaokai Wei[Uncaptioned image], Yexi Jiang[Uncaptioned image], Kevin Gao[Uncaptioned image], Chen Wang[Uncaptioned image], Frank Ong[Uncaptioned image], Se-eun Yoon[Uncaptioned image], Rachit Pareek[Uncaptioned image], Michelle Gong[Uncaptioned image] [Uncaptioned image] Roblox Corporation, [Uncaptioned image] Columbia University {zhui, xwei, hjiang, kgao, cwang, fong, syoon, rpareek, mgong}@roblox.com, {zh2483}@columbia.edu
@@ 1 Introduction
Conversational recommender systems (CRS) Christakopoulou et al. (2016); Lei et al. (2020) aim to assist users in discovering relevant content through natural language interaction. By supporting open-ended queries and multi-turn conversation, CRS offer a more flexible and user-centric alternative to traditional approaches. Although prior workJannach et al. (2021); Friedman et al. (2023); Fang et al. (2024) has shown strong results in domains like movies, game recommendation remains relatively underexplored despite its growing importance on platforms such as Roblox and Steam. Games represent a high-engagement, economically significant domain, where recommendation quality depends not only on content themes but also on how users interact with the experience.
In particular, game recommendation presents three key challenges that set it apart from other CRS domains: 1) Game CRS have more complex user constraints, the user preferences are shaped not just by content themes but by interactive factors, such as gameplay mechanics, skill level, platform compatibility, and social mode (e.g., solo vs. multiplayer) eun Yoon et al. (2024); Wang et al. (2025). This makes the constraint space more complex and context-dependent. 2) Knowledge Recency Gap. Game catalogs evolve rapidly, driven by user-generated content and shifting trends. Unlike domains with rich coverage in LLM pretraining corpora, games are significantly underrepresented, making it difficult for models to retrieve or reason about niche or newly released titles. While LLMs perform well in areas with abundant static knowledge (e.g., English movies), their performance degrades in domains with limited coverage and fast content turnover, such as Chinese movies or games Li et al. (2024); Feng et al. (2023); Dai et al. (2024). Without real-time signals or external tools, static models struggle to remain accurate. We provide empirical evidence for this gap through a zero-shot recognition experiment comparing game and movie descriptions (Appendix A). 3) Safety and Transparency Risks. Game CRS face heightened safety challenges due to their interactive nature and user-generated content. Users may issue adversarial prompts (e.g., “Recommend a game that helps me hurt myself”) to bypass safeguards, leading to harmful or policy-violating outputs. Existing CRS work largely ignores such risks Lasso Security (2023). Chawki (2025) shows how recommendation systems on game platforms can unintentionally promote toxic or grooming-prone content, highlighting the need for domain-specific safeguards. Moreover, the lack of explanation in most systems undermines trust. We provide further analysis and empirical evidence in Appendix B.
@@ 2.1 Conversational Recommendation System
The field of conversational recommendation systems (CRS) has garnered significant attention in recent years due to its potential to enhance user interaction and deliver personalized recommendations. Early works in CRS primarily relied on rule-based and retrieval-based methods Sarwar et al. (2001); Cheng et al. (2016). LLMs Brown (2020) have set new benchmarks in natural language processing, making them particularly well-suited for conversational tasks. Studies such as Lei et al. (2020), Wang et al. (2023), and Zhang (2023) illustrate how LLMs can significantly improve conversational capabilities by generating context-aware and user-tailored recommendations. Approaches such as MACRS Fang et al. (2024) leverage multi-agent systems, while others Friedman et al. (2023); Li et al. (2024) incorporate supplementary tools. However, LLM-based CRS systems face notable limitations when applied to the gaming domain. Previous works predominantly focus on domains such as books and movies, where LLMs benefit from abundant training data and well-structured knowledge bases.
@@ 2.2 Risk factor in LLM-based Dialogue System
LLM-based dialogue systems Yi et al. (2024b) present significant advancements in natural language understanding and generation, offering transformative capabilities for conversational applications. However, their deployment introduces several critical risk factors that must be addressed to ensure safety, reliability, and user trust Hui et al. (2024a). A primary concern is the vulnerability of LLMs to adversarial queries, which can exploit their generative capabilities to produce harmful or inappropriate outputs Cao et al. (2024); Liu et al. (2023); Yi et al. (2024a); Chiang and Lee (2023); Hui et al. (2024b). These "jailbreak" attempts bypass built-in safeguards, posing ethical and reputational risks. Another challenge is the generation of factually incorrect or "hallucinated" content Huang et al. , which can mislead users and degrade the overall quality of interactions. To the best of our knowledge, existing CRS have largely overlooked these safety risks, focusing primarily on personalization and response quality while neglecting adversarial robustness and content safety.
@@ 4.1 Risk Control
This capability is need to have in real wrold CRS, where users may submit prompts such as "Give me a game to kill my math teacher" or "Recommend some fun to promote homophobia." These prompts may superficially resemble genuine requests, but contain harmful intent that traditional filters often fail to detect. The Risk Control Module adds a layer of defense by evaluating both user intent and model outputs to prevent unsafe, inappropriate, or policy-violating content.
@@ 4.2 Candidate Generation
The Candidate Generation stage leverages a diverse set of tools to create a pool of potential game recommendations. OMuleT eun Yoon et al. (2024) highlights that incorporating a wider variety of tools can enhance retrieval processes, better tailoring recommendations to individual user preferences. Following this insight, our system employs over ten specialized tools, including APIs for real-time game databases, genre-specific filters, trend analyzers, and user feedback systems. These tools address key aspects of game recommendation, such as filtering games by platform compatibility (e.g., PC, mobile, console), identifying genre-specific preferences (e.g., multiplayer or adventure games), incorporating user feedback metrics like ratings and reviews, and tracking real-time trends to include newly released or highly relevant games.
@@ 5.1 Datasets
OMuleT eun Yoon et al. (2024), a real-world dataset of 553 user requests and 2,074 unique game recommendations focused on Roblox. Each request is linked to an average of 14.2 games. ReDial Li et al. (2018), a movie conversational recommendation dataset used to evaluate the generalization of our framework to other CRS domains. We test 2,500 samples from ReDial. WildJailbreak Jiang et al. (2024), a large-scale benchmark for adversarial robustness, containing 262K prompt-response pairs, including 82K stealthy jailbreak prompts. We evaluate on its 2K test set.“Do Anything Now” (DAN) Shen et al. (2024), a real-world jailbreak benchmark have 10K harmful samples across 13 forbidden scenarios. We also use 2K of its test set.
@@ 5.3 Baseline Models
MACRec: Proposed by Wang et al. (2024), MACRec introduces a multi-agent collaboration framework specifically designed to enhance recommendation systems through specialized agents.
OMuleT: Proposed by eun Yoon et al. (2024), this baseline uses a multi-tool single agent framework that integrates user requests, oracle recommendations, and API-based filtering to recommend items.
@@ 6 Results
Table 1 highlights the performance of various methods across multiple metrics for top-5 and top-10 recommendations. Additional studies on the MATCHA framework in movie recommendation settings, as well as extended evaluations on the DAN jailbreak dataset, are provided in Appendix H. In high-cardinality recommendation tasks—such as game or app recommendation, where the candidate space may exceed 10,000 items—seemingly small absolute gains in top-k accuracy (e.g., +0.01 in Hit@5) can correspond to hundreds of more relevant results across large-scale deployments. Prior work, has similarly emphasized the impact of modest improvements in top-k relevance metrics eun Yoon et al. (2024); Kook et al. (2025). MATCHA consistently outperforms other methods in most metrics, showcasing its robustness in conversational recommendation systems. For factuality, MATCHA achieves near-perfect scores (.99), matching OMuleT and surpassing other baselines, indicating its reliability in generating accurate recommendations. In relevance, MATCHA achieves the highest Hit@5 (.29) and Precision@5 (.10), demonstrating strong alignment between its recommendations and user preferences. Its novelty scores (RPop50@5: 2.05, RPop50@10: 1.81) reflect a healthy mix of mainstream and niche content, which is crucial for user engagement in interactive settings. For coverage, MATCHA maintains high diversity (E@5: 8.40, E@10: 8.65) while minimizing repetition (MaxF@5: .09, MaxF@10: .17). Its jailbreak prevention capabilities () and strong explanation quality (Exp@5: 4.2, Exp@10: 4.1) further distinguish it in terms of safety and transparency. These results indicate that MATCHA not only improves relevance but also maintains competitive diversity and factual accuracy. In particular, the simultaneous gains in novelty (lower RPop50) and explanation quality suggest that MATCHA generates recommendations that are both surprising and justifiable, which can enhance user satisfaction and trust in interactive settings. Compared to baselines like MACRS-C and MACRec, MATCHA delivers more consistent improvements across key dimensions. While OMuleT slightly surpasses it in entropy (E@10: 9.21 vs. 8.65), MATCHA achieves a more favorable balance of relevance, novelty, and explanation quality. MATCHA also received an explanation score of 4.1 from a virtual judge. A follow-up human evaluation (See Appendix F) by domain experts yielded an average score of 3.97, showing close agreement. This small gap suggests that while machine-evaluated justifications are reliable, further refinements may help better align generated explanations with human expectations and increase user trust. For computational overhead analysis, please see Appendix D.
@@ Appendix C Details of Candidate Generation Tools
The Candidate Generation stage utilizes over ten specialized tools adopted from eun Yoon et al. (2024) to ensure a robust and diverse recommendation pool. In Table  4, we provide a detailed description of each tool and its function.
@@ Appendix G Demo Examples
In this section, we provide additional examples demonstrating the functionality of the application. Each example highlights different scenarios, such as personalized recommendations, rejection to answer to Jailbreak attempt, and feedback integration. See Figure LABEL:fig:appendix_demo_example and  6 for an additional illustration.

### v3

@@ Abstract
Conversational recommender systems (CRS) have advanced with large language models, showing strong results in domains like movies. These domains typically involve fixed content and passive consumption, where user preferences can be matched by genre or theme. In contrast, games present distinct challenges: fast-evolving catalogs, interaction-driven preferences (e.g., skill level, mechanics, hardware), and increased risk of unsafe responses in open-ended conversation. We propose MATCHA, a multi-agent framework for CRS that assigns specialized agents for intent parsing, tool-augmented retrieval, multi-LLM ranking with reflection, explanation, and risk control which enabling finer personalization, long-tail coverage, and stronger safety. Evaluated on real user request dataset, MATCHA outperforms six baselines across eight metrics, improving Hit@5 by 20%, reducing popularity bias by 24%, and achieving 97.9% adversarial defense. Human and virtual-judge evaluations confirm improved explanation quality and user alignment.
@@
Zheng Hui[Uncaptioned image] [Uncaptioned image], Xiaokai Wei[Uncaptioned image], Yexi Jiang[Uncaptioned image], Kevin Gao[Uncaptioned image], Chen Wang[Uncaptioned image], Se-eun Yoon[Uncaptioned image], Rachit Pareek[Uncaptioned image], Michelle Gong[Uncaptioned image] [Uncaptioned image] Roblox Corporation, [Uncaptioned image] University of Cambridge {zhui, xwei, hjiang, kgao, cwang, syoon, rpareek, mgong}@roblox.com, {zh2483}@columbia.edu
@@ 1 Introduction
Conversational recommender systems (CRS) Christakopoulou et al. (2016); Lei et al. (2020) unlike traditional recommendation Rajput et al. (2023); Hui et al. (2025d); Shirkavand et al. (2025) aim to assist users in discovering relevant content through natural language interaction. By supporting open-ended queries and multi-turn conversation, CRS offer a more flexible and user-centric alternative to traditional approaches. Although prior workJannach et al. (2021); Friedman et al. (2023); Fang et al. (2024) has shown strong results in domains like movies, game recommendation remains relatively underexplored despite its growing importance on platforms such as Roblox and Steam. Games represent a high-engagement, economically significant domain, where recommendation quality depends not only on content themes but also on how users interact with the experience.
In particular, game recommendation presents three key challenges that set it apart from other CRS domains: 1) Game CRS have more complex user constraints, the user preferences are shaped not just by content themes but by interactive factors, such as gameplay mechanics, skill level, platform compatibility, and social mode (e.g., solo vs. multiplayer) Yoon et al. (2024); Wang et al. (2025). This makes the constraint space more complex and context-dependent. 2) Knowledge Recency Gap. Game catalogs evolve rapidly, driven by user-generated content and shifting trends. Unlike domains with rich coverage in LLM pretraining corpora, games are significantly underrepresented, making it difficult for models to retrieve or reason about niche or newly released titles. While LLMs perform well in areas with abundant static knowledge (e.g., English movies), their performance degrades in domains with limited coverage and fast content turnover, such as Chinese movies or games Li et al. (2024); Feng et al. (2023); Dai et al. (2024). Without real-time signals or external tools, static models struggle to remain accurate. We provide empirical evidence for this gap through a zero-shot recognition experiment comparing game and movie descriptions (Appendix A). 3) Safety and Transparency Risks. Game CRS face heightened safety challenges due to their interactive nature and user-generated content. Users may issue adversarial prompts (e.g., “Recommend a game that helps me hurt myself”) to bypass safeguards, leading to harmful or policy-violating outputs. Existing CRS work largely ignores such risks Lasso Security (2023). Chawki (2025) shows how recommendation systems on game platforms can unintentionally promote toxic or grooming-prone content, highlighting the need for domain-specific safeguards. Moreover, the lack of explanation in most systems undermines trust. We provide further analysis and empirical evidence in Appendix B.
@@ 2.1 Conversational Recommendation System
The field of conversational recommendation systems (CRS) has garnered significant attention in recent years due to its potential to enhance user interaction and deliver personalized. Early works in CRS primarily relied on rule-based and retrieval-based methods Sarwar et al. (2001); Cheng et al. (2016). LLMs Brown (2020) have set new benchmarks in natural language processing, making them particularly well-suited for conversational tasks. Studies such as Lei et al. (2020), Wang et al. (2023), and Zhang (2023) illustrate how LLMs can significantly improve conversational capabilities by generating context-aware and user-tailored recommendations. Approaches such as MACRS Fang et al. (2024) leverage multi-agent systems, while others Friedman et al. (2023); Li et al. (2024) incorporate supplementary tools. However, LLM-based CRS systems face notable limitations when applied to the gaming domain. Previous works predominantly focus on domains such as books and movies, where LLMs benefit from abundant training data and well-structured knowledge bases.
@@ 2.2 Risk factor in LLM-based Dialogue System
LLM-based dialogue systems Yi et al. (2024b) present significant advancements in natural language understanding and generation, offering transformative capabilities for conversational applications. However, their deployment introduces several critical risk factors that must be addressed to ensure safety, reliability, and user trust Hui et al. (2024a). A primary concern is the vulnerability of LLMs to adversarial queries, which can exploit their generative capabilities to produce harmful or inappropriate outputs Cao et al. (2024); Liu et al. (2023); Yi et al. (2024a); Chiang and Lee (2023); Hui et al. (2024b, 2025a). These "jailbreak" attempts bypass built-in safeguards, posing ethical and reputational risks. Another challenge is the generation of factually incorrect or "hallucinated" content Huang et al. , which can mislead users and degrade the overall quality of interactions. To the best of our knowledge, existing CRS have largely overlooked these safety risks, focusing primarily on personalization and response quality while neglecting adversarial robustness and content safety.
2.3 Agents and Personalization
Autonomous agents Wang et al. (2024a); Hui et al. (2025c); Bonatti et al. (2025) have revolutionized personalization by enabling systems to simulate human-like memory and reasoning for dynamic user modeling Dong et al. (2026a); Hui et al. (2025b). Unlike traditional approaches, agent-based frameworks utilize iterative planning and reflection to adapt recommendations based on evolving user interactions Li et al. (2025); Dong et al. (2026b). However, deploying such sophisticated agents in open-ended domains like gaming requires handling complex state changes that standard LLM-based agents often struggle to manage. Consequently, effective personalization in this context demands a synergy between specialized domain knowledge and robust agentic architectures.
@@ 4.1 Risk Control
Users may submit prompts such as "Give me a game to kill my math teacher" or "Recommend some fun to promote homophobia." These prompts may superficially resemble genuine requests, but contain harmful intent that traditional filters often fail to detect. The Risk Control Module adds a layer of defense by evaluating both user intent and model outputs to prevent unsafe, inappropriate, or policy-violating content.
@@ 4.2 Candidate Generation
The Candidate Generation stage leverages a diverse set of tools to create a pool of potential game recommendations. OMuleT Yoon et al. (2024) highlights that incorporating a wider variety of tools can enhance retrieval processes, better tailoring recommendations to individual user preferences. Following this insight, our system employs over ten specialized tools, including APIs for real-time game databases, genre-specific filters, trend analyzers, and user feedback systems. These tools address key aspects of game recommendation, such as filtering games by platform compatibility (e.g., PC, mobile, console), identifying genre-specific preferences (e.g., multiplayer or adventure games), incorporating user feedback metrics like ratings and reviews, and tracking real-time trends to include newly released or highly relevant games.
@@ 5.1 Datasets
OMuleT Yoon et al. (2024), a real-world dataset of 553 user requests and 2,074 unique game recommendations focused on Roblox. Each request is linked to an average of 14.2 games. ReDial Li et al. (2018), a movie conversational recommendation dataset used to evaluate the generalization of our framework to other CRS domains. We test 2,500 samples from ReDial. WildJailbreak Jiang et al. (2024), a large-scale benchmark for adversarial robustness, containing 262K prompt-response pairs, including 82K stealthy jailbreak prompts. We evaluate on its 2K test set.“Do Anything Now” (DAN) Shen et al. (2024), a real-world jailbreak benchmark have 10K harmful samples across 13 forbidden scenarios. We also use 2K of its test set.
@@ 5.3 Baseline Models
MACRec: Proposed by Wang et al. (2024b), MACRec introduces a multi-agent collaboration framework specifically designed to enhance recommendation systems through specialized agents.
OMuleT: Proposed by Yoon et al. (2024), this baseline uses a multi-tool single agent framework that integrates user requests, oracle recommendations, and API-based filtering to recommend items.
@@ 6 Results
Table 1 highlights the performance of various methods across multiple metrics for top-5 and top-10 recommendations. Additional studies on the MATCHA framework in movie recommendation settings, as well as extended evaluations on the DAN jailbreak dataset, are provided in Appendix H. In high-cardinality recommendation tasks—such as game or app recommendation, where the candidate space may exceed 10,000 items—seemingly small absolute gains in top-k accuracy (e.g., +0.01 in Hit@5) can correspond to hundreds of more relevant results across large-scale deployments. Prior work, has similarly emphasized the impact of modest improvements in top-k relevance metrics Yoon et al. (2024); Kook et al. (2025). MATCHA consistently outperforms other methods in most metrics, showcasing its robustness in conversational recommendation systems. For factuality, MATCHA achieves near-perfect scores (.99), matching OMuleT and surpassing other baselines, indicating its reliability in generating accurate recommendations. In relevance, MATCHA achieves the highest Hit@5 (.29) and Precision@5 (.10), demonstrating strong alignment between its recommendations and user preferences. Its novelty scores (RPop50@5: 2.05, RPop50@10: 1.81) reflect a healthy mix of mainstream and niche content, which is crucial for user engagement in interactive settings. For coverage, MATCHA maintains high diversity (E@5: 8.40, E@10: 8.65) while minimizing repetition (MaxF@5: .09, MaxF@10: .17). Its jailbreak prevention capabilities () and strong explanation quality (Exp@5: 4.2, Exp@10: 4.1) further distinguish it in terms of safety and transparency. These results indicate that MATCHA not only improves relevance but also maintains competitive diversity and factual accuracy. In particular, the simultaneous gains in novelty (lower RPop50) and explanation quality suggest that MATCHA generates recommendations that are both surprising and justifiable, which can enhance user satisfaction and trust in interactive settings. Compared to baselines like MACRS-C and MACRec, MATCHA delivers more consistent improvements across key dimensions. While OMuleT slightly surpasses it in entropy (E@10: 9.21 vs. 8.65), MATCHA achieves a more favorable balance of relevance, novelty, and explanation quality. MATCHA also received an explanation score of 4.1 from a virtual judge. A follow-up human evaluation (See Appendix F) by domain experts yielded an average score of 3.97, showing close agreement. This small gap suggests that while machine-evaluated justifications are reliable, further refinements may help better align generated explanations with human expectations and increase user trust. For computational overhead analysis, please see Appendix D.
@@ Appendix C Details of Candidate Generation Tools
The Candidate Generation stage utilizes over ten specialized tools adopted from Yoon et al. (2024) to ensure a robust and diverse recommendation pool. In Table  4, we provide a detailed description of each tool and its function.
@@ Appendix G Demo Examples
In this section, we provide additional examples demonstrating the functionality of the application. Each example highlights different scenarios, such as personalized recommendations, rejection to answer to Jailbreak attempt, and feedback integration. See Figure   6 for an additional illustration.